if (a+2)x^2- 2ax - a = 0

what is range of a ?

Answer by AnlytcPhil(1276)


if (a+2)x^2- 2ax - a = 0
`
what is range of a ?
`
If you were to solve this problem by the quadratic formula,
`
A = (a+2), B = -2a, C = -a
`
the expression underneath the square root, B2 - 4AC, known as the
'discriminant', must not be negative.
`
` ` ` ` ` `B2 - 4AC >= 0
`
(-2a)2 - 4(a+2)(-a) >= 0
`
` ` ` 4a2 + 4a(a+2) >= 0
`
` ` `4a2 + 4a2 + 8a >= 0
`
` ` ` ` ` `8a2 + 8a >= 0
`
` ` ` ` ` 8a(a + 1) >= 0
`
The zeros of the left side are 0 and -1
`
Make a number line, and mark these solid since they are solutions.
`
----------•--•------------
-4 -3 -2 -1 `0 +1 +2 +3 +4
`
Choose any value in the leftmost region, the region left of -1. The easiest
value to choose is -2. Substitute it into the expression
`
8a(a + 1)
8(-2)(-2 + 1) = -16(-1) = +16. This is nonnegative, so shade all those values.
`
<==========•--•------------
`-4 -3 -2 -1 0 +1 +2 +3 +4
`
Choose any value in the middle region, the region between -1 and 0. The
easiest value to choose is -1/2. Substitute it into the expression
`
8a(a + 1)
8(-1/2)(-1/2 + 1) = -4(1/2) = -2. This is negative, so do not shade the values
in the middle region.
`
Choose any value in the rightmost region, the region right of 0. The easiest
value to choose is 1. Substitute it into the expression
`
8a(a + 1)
8(1)(1 + 1) = 8(2) = +16. This is nonnegative, so shade all those values.
`
<==========•--•============>
`-4 -3 -2 -1 `0 +1 +2 +3 +4
`
The interval notation for this is
`
(-¥,-1] U [0, ¥)
`
Edwin
AnlytcPhil@aol.com